OpenStudy (anonymous):

Solve the equation for y. 2y - 4/3 = y + 9/2 can someone please help me find y

OpenStudy (anonymous):

2y - 4/3 = y + 9/2 2y - 4/3 + 4/3 = y + 9/2 + 4/3 2y = y + 27/9 + 12/9 2y - y = y - y + 39/9 y = 39/9 = 13/3

OpenStudy (anonymous):

thank u but can u explan how u got that please